# Bilinear spherical maximal function with fractal dilations

### v1

@@ 5. Estimate for local maximal operator
$\leq n_{2}n_{1}^{\frac{1}{3}}2^{-|\mathbf{n}|\frac{d-3-\beta}{3}}\bigg(\sum_{k\in{\mathbb{Z}}}\sum_{i\in{\mathbb{Z}}}\min\Big\{1,2^{\frac{2(i-k)}{3}},2^{\frac{2(k+n_{1}-i)}{3}}\Big\}\|h^{i}_{1}\|_{L^{1}}\bigg)^{\frac{2}{3}}\bigg(\sum_{k\in{\mathbb{Z}}}\|\mathbf{R}_{k+n_{2}}f_{2}\|^{2}_{L^{2}}\bigg)^{\frac{1}{3}}$
$\lesssim n_{2}n_{1}2^{-|\mathbf{n}|\frac{d-3-\beta}{3}}\bigg(\sum_{i\in{\mathbb{Z}}}\|h^{i}_{1}\|_{L^{1}}\bigg)^{\frac{2}{3}}\|f_{2}\|_{L^{2}}^{\frac{2}{3}}$
$\lesssim|\mathbf{n}|^{2}2^{-|\mathbf{n}|\frac{d-3-\beta}{3}},$
@@
This completes the proof of part $(1)$ of Lemma 6.2 under the assumption that Lemma 6.3 holds.
@@
We apply part $(2)$ of Lemma 5.2 to get that
$\int_{\mathbb{R}^{d}\setminus\mathcal{F}}\Big|\mathcal{M}^{\mathbf{n}}_{k}(h^{i}_{1},f_{2})(x)\Big|^{\frac{2}{3}}dx$ $\lesssim$ $2^{-|\mathbf{n}|\frac{d-3-\beta}{3}}\|\mathbf{R}_{k+n_{1}}h^{i}_{1}\|^{\frac{2}{3}}_{L^{1}}\|\mathbf{R}_{k+n_{2}}f_{2}\|^{\frac{2}{3}}_{L^{2}}$
$\lesssim$ $2^{-|\mathbf{n}|\frac{d-3-\beta}{3}}\|h^{i}_{1}\|^{\frac{2}{3}}_{L^{1}}\|f_{2}\|^{\frac{2}{3}}_{L^{2}}.$
@@
$\|\mathcal{M}^{(m,n_{2})}_{k}(h^{i}_{1}\ast\varphi_{2^{-k-n_{1}}}^{i},f_{2})\|_{L^{2/3}}\lesssim 2^{-|(m,n_{2})|\frac{d-3-\beta}{2}}\|h^{i}_{1}\ast\varphi_{2^{-k-n_{1}}}^{i}\|_{L^{1}}\|f_{2}\|_{L^{2}}.$ (6.3)
Let us assume the estimate (6.3) for a moment. Note that $\psi\in\mathcal{S}({\mathbb{R}}^{d})$, we have $|\psi(w)|\lesssim_{N}(1+|w|)^{-N}$ for every $N>1$. Consequently, for $N\geq d+1$, we get that
@@
$\lesssim\sum_{m\geq 0}2^{-|(m,n_{2})|\frac{d-3-\beta}{3}}\|h^{i}_{1}\ast\varphi_{2^{-k-n_{1}}}^{i}\|_{L^{1}}^{\frac{2}{3}}\|f_{2}\|_{L^{2}}^{\frac{2}{3}}$
$\lesssim\sum_{m\geq 0}2^{-|(m,n_{2})|\frac{d-3-\beta}{3}}2^{-\frac{2n_{1}d}{3}}2^{\frac{2(i-k)}{3}}\|h^{i}_{1}\|_{L^{1}}^{\frac{2}{3}}\|f_{2}\|_{L^{2}}^{\frac{2}{3}}$
$\leq 2^{-\frac{2n_{1}d}{3}}2^{\frac{2(i-k)}{3}}\bigg(\sum_{m=0}^{n_{2}-1}2^{-n_{2}\frac{d-3-\beta}{3}}\|h^{i}_{1}\|_{L^{1}}^{\frac{2}{3}}\|f_{2}\|_{L^{2}}^{\frac{2}{3}}+\sum_{m\geq n_{2}}2^{-m\frac{d-3-\beta}{3}}\|h^{i}_{1}\|_{L^{1}}^{\frac{2}{3}}\|f_{2}\|_{L^{2}}^{\frac{2}{3}}\bigg)$
$\leq n_{2}2^{-\frac{2n_{1}d}{3}}2^{\frac{2(i-k)}{3}}2^{-n_{2}\frac{d-3-\beta}{3}}\|h^{i}_{1}\|_{L^{1}}^{\frac{2}{3}}\|f_{2}\|_{L^{2}}^{\frac{2}{3}}$
$\leq n_{2}2^{-|\mathbf{n}|\frac{d-3-\beta}{3}}2^{\frac{2(i-k)}{3}}\|h^{i}_{1}\|_{L^{1}}^{\frac{2}{3}}\|f_{2}\|_{L^{2}}^{\frac{2}{3}}.$
@@
When $m\geq 1$, the desired estimate in (6.3) follows by adapting the proof of Lemma 5.2 to the scale $k$. Thus, it remains to prove (6.3) for $m=0$.
Let $b_{1}=h^{i}_{1}\ast\varphi_{2^{-k-n_{1}}}^{i}$. In view of the discussion in Section 4, we need to prove the desired $L^{p}$-boundedness for the maximal operator corresponding to bilinear operators defined by
$T^{n_{2}}_{t,k}(b_{1},f_{2})(x)=\int_{{\mathbb{R}}^{d}\times{\mathbb{R}}^{d}}\widehat{P_{k}b_{1}}(\xi)\widehat{R_{k+n_{2}}f_{2}}(\eta)m(t\xi,t\eta)e^{2\pi it|(\xi,\eta)|}e^{2\pi ix\cdot(\xi,\eta)}\;d\xi d\eta,\;t\in E_{k},$
where $m(\xi,\eta)=|(\xi,\eta)|^{-\frac{2d-1}{2}}$. Again, we restrict ourselves to the case of unit scale, that is $k=0$. The following arguments can be easily adapted to obtain the corresponding estimates for a general scale $k$. As previously, we denote $T^{n_{2}}_{t,0}$ by $T^{n_{2}}_{t}$.
Let $N$ be a positive integer such that $N\sim 2^{n_{2}}$ and write
$e^{2\pi it|(\xi,\eta)|}=e^{2\pi it\Phi_{N}(\xi,\eta)}e^{2\pi it\sqrt{|\xi|^{2}+N^{2}}},$
@@
$\Phi_{N}(\xi,\eta):=|(\xi,\eta)|-\sqrt{|\xi|^{2}+N^{2}}=\frac{(|\eta|-N)(|\eta|+N)}{\sqrt{|\xi|^{2}+|\eta|^{2}}+\sqrt{|\xi|^{2}+N^{2}}}.$
Note that $\Phi_{N}$ satisfies the following estimate
$|\partial^{\alpha}_{\xi}\Phi_{N}(\xi,\eta)|\lesssim 1,$ (6.4)
for all multi-indices $\alpha$ uniformly in $(\xi,\eta)$ when $|\xi|\leq 2$ and $N\leq|\eta|<N+1$. This gives us the following decomposition of the operator $T^{n_{2}}_{t}.$
$T^{n_{2}}_{t}(b_{1},f_{2})(x)$ $=\sum_{N=2^{n_{1}-1}}^{2^{n_{1}+1}}T^{n_{2}}_{t,N}(b_{1},f_{2})(x),$
@@
$T^{n_{2}}_{t,N}(b_{1},f_{2})(x)$ $=\int_{{\mathbb{R}}^{d}\times{\mathbb{R}}^{d}}m_{t}^{n_{2}}(\xi,\eta)e^{2\pi it\Phi_{N}(\xi,\eta)}\widehat{U_{t,N}b_{1}}(\xi)\widehat{F_{N}f_{2}}(\eta)e^{2\pi ix\cdot(\xi,\eta)}\;d\xi d\eta,$
$\widehat{U_{t,N}b_{1}}(\xi):=\widehat{b_{1}}(\xi)\widehat{\phi}\left(\xi\right)e^{2\pi it\left(\sqrt{|\xi|^{2}+N^{2}}\right)},$ (6.5)
@@
$m^{n_{2}}_{t}(\xi,\eta):=m(t\xi,t\eta)\widehat{\phi}\left(\xi\right)\widehat{\psi}\left(2^{-n_{2}}\eta\right).$
@@
$|\partial^{\alpha}_{\xi}\big(m^{n_{2}}_{t}(\xi,\eta)e^{2\pi it\Phi_{N}(\xi,\eta)}\big)|\lesssim C_{\alpha}2^{-\frac{2d-1}{2}|n_{2}|},$ (6.6)
@@
Since $\widehat{U_{t,N}b_{1}}(\eta)=\int_{{\mathbb{R}}^{d}}U_{t,N}b_{1}(y)e^{-2\pi i\xi\cdot y}~dy$, we can rewrite
$T^{n_{2}}_{t,N}(b_{1},f_{2})(x)$ $=\int_{{\mathbb{R}}^{d}\times{\mathbb{R}}^{d}}\widehat{F_{N}f_{2}}(\eta)U_{t,N}b_{1}(y)\Big(\int_{{\mathbb{R}}^{d}}m^{n_{2}}_{t}(\xi,\eta)e^{2\pi it\Phi_{N}(\xi,\eta)}e^{2\pi i\xi\cdot(x-y)}d\xi\Big)e^{2\pi ix\cdot\eta}\;d\eta dy.$ (6.7)
@@
$\int_{{\mathbb{R}}^{d}}m^{n_{2}}_{t}(\xi,\eta)e^{2\pi it\Phi_{N}(\xi,\eta)}e^{2\pi i\xi\cdot(x-y)}d\xi$ (6.8)
$=\frac{1}{(1+4\pi^{2}|x-y|^{2})^{\lceil\frac{d}{2}\rceil+1}}\int_{{\mathbb{R}}^{d}}(1-\Delta_{\xi})^{\lceil\frac{d}{2}\rceil+1}\Big(m^{n_{2}}_{t}(\xi,\eta)e^{2\pi it\Phi_{N}(\xi,\eta)}\Big)e^{2\pi i\xi\cdot(x-y)}d\xi,$
@@
$B^{n_{2}}_{t,N}\left(F_{N}f_{2}\right)(x,\xi):=$ $\int_{\mathbb{R}^{d}}\left(1-\Delta_{\xi}\right)^{\lceil\frac{d}{2}\rceil+1}\left(m^{n_{2}}_{t}(\xi,\eta)e^{2\pi it\Phi_{N}(\xi,\eta)}\right)\widehat{F_{N}f_{2}}(\eta)e^{2\pi ix\cdot\eta}d\eta,$ (6.9)
@@
$A\left(U_{t,N}b_{1}\right)(x,\eta):=\int_{\mathbb{R}^{d}}\frac{U_{t,N}b_{1}(y)e^{2\pi i\xi\cdot(x-y)}}{\left(1+4\pi^{2}\left|x-y\right|^{2}\right)^{\lceil\frac{d}{2}\rceil+1}}dy.$
@@
$|A\left(U_{t,N}b_{1}\right)(x,\eta)|$ $\leq\int_{\mathbb{R}^{d}}\frac{|U_{t,N}b_{1}(y)|}{\left(1+4\pi^{2}\left|x-y\right|^{2}\right)^{\lceil\frac{d}{2}\rceil+1}}dy$
@@
This gives us the following bound.
$|T^{n_{2}}_{t,N}(b_{1},f_{2})(x)|\leq L\left(U_{t,N}b_{1}\right)(x)\int_{\mathbb{R}^{d}}\left|B^{n_{2}}_{t,N}\left(F_{N}f_{2}\right)(x,\xi)\right|d\xi.$ (6.10)
With this decomposition in mind, we follow the arguments from the proof of estimate (5.1) to get that
$\mathcal{M}^{(0,n_{2})}_{0}(b_{1},f_{2})(x)\lesssim$ $\sum_{N=2^{n_{2}-1}}^{2^{n_{2}+1}}2^{\frac{n_{2}}{2}}\mathcal{G}^{n_{2}}_{0,N}(b_{1},f_{2})(x)+\sum_{N=2^{n_{2}-1}}^{2^{n_{2}+1}}\big(\mathcal{G}^{n_{2}}_{0,N}(b_{1},f_{2})(x)\tilde{\mathcal{G}}^{n_{2}}_{0,N}(b_{1},f_{2})(x)\big)^{\frac{1}{2}},$
where
$\mathcal{G}^{n_{2}}_{0,N}(b_{1},f_{2})(x):=\left(\sum_{I\in\mathcal{E}_{0}}\int_{I}|\mathcal{A}_{s}(P_{0}b_{1},R_{n_{2}}F_{N}f_{2})(x)|^{2}\;ds\right)^{\frac{1}{2}},$
$\tilde{\mathcal{G}}^{n_{2}}_{0,N}(b_{1},f_{2})(x):=\left(\sum_{I\in\mathcal{E}_{0}}\int_{I}\left|\frac{\partial}{\partial s}\mathcal{A}_{s}(P_{0}b_{1},R_{n_{2}}F_{N}f_{2})(x)\right|^{2}\;ds\right)^{\frac{1}{2}},$
and $\mathcal{E}_{0}$ is a collection of disjoint intervals of length $2^{-n_{2}}$ and covering $E_{0}$.
Since the multiplier $2^{-|\mathbf{n}|}\frac{\partial}{\partial s}\widehat{\sigma}(s\xi,s\eta)$ behaves like $\widehat{\sigma}(s\xi,s\eta)$ where $1\leq s\leq 2$. Therefore, it is enough to prove the desired estimates for the square function $\Big(\sum\limits_{I\in\mathcal{E}_{0}}\int_{I}|\mathcal{A}_{s}(P_{0}f_{1},R_{n_{2}}F_{N}f_{2})(x)|^{2}\;ds\Big)^{\frac{1}{2}}$.
We invoke the decomposition of the multiplier in (6.10) to get that
$\mathcal{G}^{n_{2}}_{0,N}(b_{1},f_{2})(x)$
$\lesssim\left(\int_{\mathbb{R}^{d}}\sup_{1\leq s\leq 2}\left|B^{n_{2}}_{s,N}\left(F_{N}f_{2}\right)(x,\xi)\right|d\xi\right)\bigg(\sum_{I\in\mathcal{E}_{0}}\int_{I}\left|L\left(U_{s,N}b_{1}\right)(x)\right|^{2}\;ds\bigg)^{\frac{1}{2}}+error~terms.$
Note that for the error terms, we have better decay of the multiplier, see (4.2). Thus, it is enough to prove desired bounds for the leading term.
Recall that $\operatorname{supp}(\widehat{\phi})\subset B(0,2)$. Similar to the $L^{p}$-estimates in Lemma 4.1, we can get that
$\left\|\int_{\mathbb{R}^{d}}\sup_{1\leq s\leq 2}\left|B^{n_{2}}_{s,N}\left(F_{N}f_{2}\right)(\cdot,\xi)\right|d\xi\right\|_{L^{2}}\lesssim 2^{-n_{2}\frac{2d-1}{2}}\|f_{2}\|_{L^{2}}.$

### v2

@@ 5. Estimate for local maximal operator $\mathcal{M}_{E_{0}}$
$\leq n_{2}n_{1}^{\frac{1}{3}}\min\left\{2^{-|\mathbf{n}|\frac{d-2-3\beta}{3}},2^{-|\mathbf{n}|\frac{d-3-\beta}{3}}\right\}\bigg(\sum_{k\in{\mathbb{Z}}}\sum_{i\in{\mathbb{Z}}}\min\Big\{1,2^{\frac{2(i-k)}{3}},2^{\frac{2(k+n_{1}-i)}{3}}\Big\}\|h^{i}_{1}\|_{L^{1}}\bigg)^{\frac{2}{3}}$
$\hskip 113.81102pt\times\bigg(\sum_{k\in{\mathbb{Z}}}\|\mathbf{R}_{k+n_{2}}f_{2}\|^{2}_{L^{2}}\bigg)^{\frac{1}{3}}$
$\lesssim n_{2}n_{1}\min\left\{2^{-|\mathbf{n}|\frac{d-2-3\beta}{3}},2^{-|\mathbf{n}|\frac{d-3-\beta}{3}}\right\}\bigg(\sum_{i\in{\mathbb{Z}}}\|h^{i}_{1}\|_{L^{1}}\bigg)^{\frac{2}{3}}\|f_{2}\|_{L^{2}}^{\frac{2}{3}}$
$\lesssim|\mathbf{n}|^{2}\min\left\{2^{-|\mathbf{n}|\frac{d-2-3\beta}{3}},2^{-|\mathbf{n}|\frac{d-3-\beta}{3}}\right\},$
@@
This proves (6.2) and completes the proof of part $(1)$ of Lemma 6.2 under the assumption that Lemma 6.4 holds. ∎
@@
Using a scaling argument along with part $(2)$ of Lemma 5.1, we get
$\int_{\mathbb{R}^{d}\setminus\mathcal{F}}\Big|\mathcal{M}^{\mathbf{n}}_{k}(h^{i}_{1},f_{2})(x)\Big|^{\frac{2}{3}}dx$ $\lesssim\min\left\{2^{-|\mathbf{n}|\frac{d-2-3\beta}{3}},2^{-|\mathbf{n}|\frac{d-3-\beta}{3}}\right\}\|\mathbf{R}_{k+n_{1}}h^{i}_{1}\|^{\frac{2}{3}}_{L^{1}}\|\mathbf{R}_{k+n_{2}}f_{2}\|^{\frac{2}{3}}_{L^{2}}$ (6.3)
$\lesssim\min\left\{2^{-|\mathbf{n}|\frac{d-2-3\beta}{3}},2^{-|\mathbf{n}|\frac{d-3-\beta}{3}}\right\}\|h^{i}_{1}\|^{\frac{2}{3}}_{L^{1}}\|f_{2}\|^{\frac{2}{3}}_{L^{2}}.$
@@
$\|\mathcal{M}^{(m,n_{2})}_{k}(h^{i}_{1}\ast\varphi_{2^{-k-n_{1}}}^{i},f_{2})\|_{L^{2/3}}\lesssim\min\left\{2^{-|(m,n_{2})|\frac{d-2-3\beta}{2}},2^{-|(m,n_{2})|\frac{d-3-\beta}{2}}\right\}\|h^{i}_{1}\ast\varphi_{2^{-k-n_{1}}}^{i}\|_{L^{1}}\|f_{2}\|_{L^{2}}.$ (6.4)
Let us assume the estimate (6.4) for a moment. Since $\psi\in\mathcal{S}({\mathbb{R}}^{d})$, we have $|\psi(w)|\lesssim_{N}(1+|w|)^{-N}$ for every $N>1$. Consequently, for $N\geq d+1$, we get that
@@
$\lesssim\sum_{m\geq 0}\min\left\{2^{-|(m,n_{2})|\frac{d-2-3\beta}{3}},2^{-|(m,n_{2})|\frac{d-3-\beta}{3}}\right\}\|h^{i}_{1}\ast\varphi_{2^{-k-n_{1}}}^{i}\|_{L^{1}}^{\frac{2}{3}}\|f_{2}\|_{L^{2}}^{\frac{2}{3}}$
$\lesssim\sum_{m\geq 0}\min\left\{2^{-|(m,n_{2})|\frac{d-2-3\beta}{3}},2^{-|(m,n_{2})|\frac{d-3-\beta}{3}}\right\}2^{-\frac{2n_{1}d}{3}}2^{\frac{2(i-k)}{3}}\|h^{i}_{1}\|_{L^{1}}^{\frac{2}{3}}\|f_{2}\|_{L^{2}}^{\frac{2}{3}}$
$\leq 2^{-\frac{2n_{1}d}{3}}2^{\frac{2(i-k)}{3}}\bigg(\sum_{m=0}^{n_{2}-1}\min\left\{2^{-n_{2}\frac{d-2-3\beta}{3}},2^{-n_{2}\frac{d-3-\beta}{3}}\right\}\|h^{i}_{1}\|_{L^{1}}^{\frac{2}{3}}\|f_{2}\|_{L^{2}}^{\frac{2}{3}}$
$\hskip 85.35826pt+\sum_{m\geq n_{2}}\min\left\{2^{-m\frac{d-2-3\beta}{3}},2^{-m\frac{d-3-\beta}{3}}\right\}\|h^{i}_{1}\|_{L^{1}}^{\frac{2}{3}}\|f_{2}\|_{L^{2}}^{\frac{2}{3}}\bigg)$
$\leq n_{2}2^{-\frac{2n_{1}d}{3}}2^{\frac{2(i-k)}{3}}\min\left\{2^{-n_{2}\frac{d-2-3\beta}{3}},2^{-n_{2}\frac{d-3-\beta}{3}}\right\}\|h^{i}_{1}\|_{L^{1}}^{\frac{2}{3}}\|f_{2}\|_{L^{2}}^{\frac{2}{3}}$
$\leq n_{2}\min\left\{2^{-|\mathbf{n}|\frac{d-2-3\beta}{3}},2^{-|\mathbf{n}|\frac{d-3-\beta}{3}}\right\}2^{\frac{2(i-k)}{3}}\|h^{i}_{1}\|_{L^{1}}^{\frac{2}{3}}\|f_{2}\|_{L^{2}}^{\frac{2}{3}}.$
@@
Due to the scaling argument in (6.1), it is enough to prove (6.4) for $\mathcal{M}^{(m,n_{2})}_{0}$. Hence, when $m\geq 1$, the desired estimate in (6.4) follows from Lemma 5.1. Thus, it remains to prove (6.4) for $m=0$.
Let $b_{1}=h^{i}_{1}\ast\varphi_{2^{-k-n_{1}}}^{i}$ and denote $\mathcal{M}^{(0,n_{2})}_{0}=\mathcal{M}^{n_{2}}_{0}$. Performing a decomposition similar to that in (5.1), we have
$\mathcal{M}^{n_{2}}_{0}(b_{1},f_{2})(x)\leq\mathbb{I}^{n_{2}}(x)+\mathbb{II}^{n_{2}}(x)+\mathbb{III}^{n_{2}}(x),$
@@
$\mathbb{I}^{n_{2}}(x)=\sup_{t\in E_{0}}\left|\int_{{\mathbb{R}}^{d}\times{\mathbb{R}}^{d}}\frac{a_{0}e^{it|(\xi,\eta)|}}{(t|(\xi,\eta)|)^{\frac{2d-1}{2}}}\widehat{\phi}\left(\xi\right)\widehat{\psi}\left(2^{-n_{2}}\eta\right)\widehat{b_{1}}(\xi)\widehat{f_{2}}(\eta)e^{ix\cdot(\xi+\eta)}~d\xi d\eta\right|,$
$\mathbb{II}^{n_{2}}(x)=\sup_{t\in E_{0}}\left|\int_{{\mathbb{R}}^{d}\times{\mathbb{R}}^{d}}\frac{b_{0}e^{-it|(\xi,\eta)|}}{(t|(\xi,\eta)|)^{\frac{2d-1}{2}}}\widehat{\phi}\left(\xi\right)\widehat{\psi}\left(2^{-n_{2}}\eta\right)\widehat{b_{1}}(\xi)\widehat{f_{2}}(\eta)e^{ix\cdot(\xi+\eta)}~d\xi d\eta\right|,$
and
$\mathbb{III}^{n_{2}}(x)=\sup_{t\in E_{0}}\left|\int_{{\mathbb{R}}^{d}\times{\mathbb{R}}^{d}}\mathfrak{m}_{t}^{n_{2}}(\xi,\eta)\widehat{b_{1}}(\xi)\widehat{f_{2}}(\eta)e^{ix\cdot(\xi+\eta)}~d\xi d\eta\right|.$
Here $\mathfrak{m}_{t}^{n_{2}}(\xi,\eta)=\Big[\widehat{d\sigma}(t\xi,t\eta)-\Big(\frac{a_{0}e^{it|(\xi,\eta)|}}{(t|(\xi,\eta)|)^{\frac{2d-1}{2}}}+\frac{b_{0}e^{-it|(\xi,\eta)|}}{(t|(\xi,\eta)|)^{\frac{2d-1}{2}}}\Big)\Big]\widehat{\phi}\left(\xi\right)\widehat{\psi}\left(2^{-n_{2}}\eta\right)$. Using the estimate (4.2) with $M=0$, and the support condition of $\widehat{\phi}$ and $\widehat{\psi}$, we have
$\big|\partial^{\alpha}_{\xi}\partial^{\beta}_{\eta}\mathfrak{m}_{t}^{n_{2}}(\xi,\eta)\big|\lesssim 2^{-n_{2}\frac{2d-1}{2}}|\eta|^{-|\beta|},\text{ for all }t\in E_{0}.$
This yields that
$\mathbb{III}^{n_{2}}(x)\lesssim 2^{-n_{2}\frac{2d-1}{2}}M_{HL}f_{1}(x)M_{HL}f_{2}(x).$
Therefore, $\mathbb{III}^{n_{2}}$ satisfies the desired bounds in (6.4).
Since, the arguments of proof to address the terms $\mathbb{I}^{n_{2}}$ and $\mathbb{II}^{n_{2}}$ are similar, we provide the proof of (6.4) for the term $\mathbb{I}^{n_{2}}$.
Let $\mathcal{E}_{0}$ be the disjoint collection of intervals of length $2^{-n_{2}}$ covering $E_{0}$. Decomposing $\mathbb{I}^{n_{2}}$ similar to that of $\mathbb{I}^{\mathbf{n}}$ in Section 5.1, we have
$\mathbb{I}^{n_{2}}(x)$ $\lesssim\sum_{N=2^{n_{2}-1}}^{2^{n_{2}+1}}\sup_{I\in\mathcal{E}_{0}}\sup_{t\in I}\Big|T^{n_{2}}_{t,N}\Big(b_{1},f_{2}\Big)(x)\Big|,$
@@
$T^{n_{2}}_{t,N}\Big(b_{1},f_{2}\Big)(x):=\int_{{\mathbb{R}}^{d}\times{\mathbb{R}}^{d}}\frac{e^{it|(\xi,\eta)|}}{(t|(\xi,\eta)|)^{\frac{2d-1}{2}}}\widehat{\phi}\left(\xi\right)\widehat{\psi}\left(2^{-n_{2}}\eta\right)\widehat{b_{1}}(\xi)\widehat{F_{N}f_{2}}(\eta)e^{ix\cdot(\xi+\eta)}~d\xi d\eta,$
and $F_{N}$ is defined as in (4.4).
Then, from (5.3) and $\ell^{1}\hookrightarrow\ell^{\infty}$ embedding for $\mathcal{E}_{0}$, we get
$\mathbb{I}^{n_{2}}(x)\lesssim$ $\sum_{N=2^{n_{2}-1}}^{2^{n_{2}+1}}\sum_{I\in\mathcal{E}_{0}}2^{\frac{n_{2}}{2}}\mathcal{G}^{n_{2}}_{I,N}(b_{1},f_{2})(x)+\sum_{N=2^{n_{2}-1}}^{2^{n_{2}+1}}\sum_{I\in\mathcal{E}_{0}}\big(\mathcal{G}^{n_{2}}_{I,N}(b_{1},f_{2})(x)\tilde{\mathcal{G}}^{n_{2}}_{I,N}(b_{1},f_{2})(x)\big)^{\frac{1}{2}},$
where
$\mathcal{G}^{n_{2}}_{I,N}(b_{1},f_{2})(x):=\left(\sum_{I\in\mathcal{E}_{0}}\int_{I}|T^{n_{2}}_{t,N}(b_{1},f_{2})(x)|^{2}\;ds\right)^{\frac{1}{2}}~\text{ and}$
$\tilde{\mathcal{G}}^{n_{2}}_{I,N}(b_{1},f_{2})(x):=\left(\sum_{I\in\mathcal{E}_{0}}\int_{I}\left|\frac{\partial}{\partial s}T^{n_{2}}_{t,N}(b_{1},f_{2})(x)\right|^{2}\;ds\right)^{\frac{1}{2}}.$
Fix a positive integer NN such that $2^{n_{2}-1}\leq N\leq 2^{n_{2}+1}$ and write
$e^{it|\xi,\eta)|}=e^{it\tilde{\Phi}(\xi,\eta)}e^{it\sqrt{|\xi|^{2}+N^{2}}},$
where
$\tilde{\Phi}_{N}(\xi,\eta):=|(\xi,\eta)|-\sqrt{|\xi|^{2}+N^{2}}=\frac{(|\eta|-N)(|\eta|+N)}{\sqrt{|\xi|^{2}+|\eta|^{2}}+\sqrt{|\xi|^{2}+N^{2}}}.$
Note that due to support of $\widehat{\phi}$ and $\widehat{\psi}$, we have $|\xi|\leq 2$ and $2^{n_{2}-1}\leq|\eta|\leq 2^{n_{2}+1},n_{2}\geq 1$. Thus, $\tilde{\Phi}_{N}$ satisfies
$|\partial^{\alpha}_{\xi}\tilde{\Phi}_{N}(\xi,\eta)|\lesssim 1,$ (6.5)
for all multi-indices $\alpha$. Then, we can write
$T^{n_{2}}_{t,N}(b_{1},f_{2})(x)$ $=\int_{{\mathbb{R}}^{d}\times{\mathbb{R}}^{d}}m_{t}^{n_{2}}(\xi,\eta)\widehat{U_{t,N}b_{1}}(\xi)\widehat{F_{N}f_{2}}(\eta)e^{ix\cdot(\xi+\eta)}\;d\xi d\eta,$
where
$\widehat{U_{t,N}b_{1}}(\xi):=\widehat{b_{1}}(\xi)\widehat{\phi}\left(\xi\right)e^{it\left(\sqrt{|\xi|^{2}+N^{2}}\right)},$
@@
$m^{n_{2}}_{t}(\xi,\eta):=\frac{e^{it\tilde{\Phi}_{N}(\xi,\eta)}}{|t(\xi,\eta)|}\widehat{\phi}\left(\xi\right)\widehat{\psi}\left(2^{-n_{2}}\eta\right).$
@@
$|\partial^{\alpha}_{\xi}\big(m^{n_{2}}_{t}(\xi,\eta)\big)|\lesssim C_{\alpha}2^{-\frac{2d-1}{2}n_{2}},$ (6.6)
@@
Since $\widehat{U_{t,N}b_{1}}(\eta)=\int_{{\mathbb{R}}^{d}}U_{t,N}b_{1}(y)e^{-i\xi\cdot y}~dy$, we can rewrite
$T^{n_{2}}_{t,N}(b_{1},f_{2})(x)$ $=\int_{{\mathbb{R}}^{d}\times{\mathbb{R}}^{d}}\widehat{F_{N}f_{2}}(\eta)U_{t,N}b_{1}(y)\Big(\int_{{\mathbb{R}}^{d}}m^{n_{2}}_{t}(\xi,\eta)e^{i\xi\cdot(x-y)}d\xi\Big)e^{ix\cdot\eta}\;d\eta dy.$ (6.7)
@@
$\int_{{\mathbb{R}}^{d}}m^{n_{2}}_{t}(\xi,\eta)e^{i\xi\cdot(x-y)}d\xi=\frac{1}{(1+|x-y|^{2})^{\lceil\frac{d}{2}\rceil+1}}\int_{{\mathbb{R}}^{d}}(1-\Delta_{\xi})^{\lceil\frac{d}{2}\rceil+1}\Big(m^{n_{2}}_{t}(\xi,\eta)\Big)e^{i\xi\cdot(x-y)}d\xi.$ (6.8)
@@
$B^{n_{2}}_{t,N}\left(F_{N}f_{2}\right)(x,\xi):=$ $\int_{\mathbb{R}^{d}}\left(1-\Delta_{\xi}\right)^{\lceil\frac{d}{2}\rceil+1}\left(m^{n_{2}}_{t}(\xi,\eta)\right)\widehat{F_{N}f_{2}}(\eta)e^{ix\cdot\eta}d\eta,$
@@
$A\left(U_{t,N}b_{1}\right)(x,\eta):=\int_{\mathbb{R}^{d}}\frac{U_{t,N}b_{1}(y)e^{i\xi\cdot(x-y)}}{\left(1+\left|x-y\right|^{2}\right)^{\lceil\frac{d}{2}\rceil+1}}dy.$
@@
$|A\left(U_{t,N}b_{1}\right)(x,\eta)|$ $\leq\int_{\mathbb{R}^{d}}\frac{|U_{t,N}b_{1}(y)|}{\left(1+\left|x-y\right|^{2}\right)^{\lceil\frac{d}{2}\rceil+1}}dy$
@@
Then, we have
$|T^{n_{2}}_{t,N}(b_{1},f_{2})(x)|\leq L\left(U_{t,N}b_{1}\right)(x)\int_{\mathbb{R}^{d}}\left|B^{n_{2}}_{t,N}\left(F_{N}f_{2}\right)(x,\xi)\right|d\xi,$
and
$\mathcal{G}^{n_{2}}_{I,N}(b_{1},f_{2})(x)\lesssim\left(\int_{\mathbb{R}^{d}}\sup_{s\in I}\left|B^{n_{2}}_{s,N}\left(f_{2}\right)(x,\eta)\right|d\eta\right)\bigg(\int_{I}\left|L\left(U_{s,N}^{n_{2}}b_{1}\right)(x)\right|^{2}\;ds\bigg)^{\frac{1}{2}}.$
Recall that $\operatorname{supp}(\widehat{\phi})\subset B(0,2)$ and $|I|=2^{-n_{2}}$. Similarly to the $L^{p}$-estimates in Lemma 4.1 along with (6.6), we can get that
$\int_{\mathbb{R}^{d}}\left\|\sup_{s\in I}\left|B^{n_{2}}_{s,N}\left(F_{N}f_{2}\right)(\cdot,\xi)\right|\right\|_{L^{2}}d\xi\lesssim 2^{-n_{2}d}\|f_{2}\|_{L^{2}}.$